Assume that Morton Lumber and Building offers a 2/10, n30 discount . Further assume that we did not take advantage of this discount. What would the Effective Annual Percentage Rate be because we did not pay within 10 days and get the 2% discount?

Question 1 options:

10.39

15.12

6.79

44.56

Pleas explain the difference of taking advantage of or foregoing the cash discount..
Answer #1

The effective annual rate is calculated by using one simple formula given below:

Discount is 2% when payable within 10 days out of 30 days

Step 1

30-10= 20

Step 2

365/20

= 18.25

Step 3

Substracting 2% from 1

= 1-0.0= 98

= 1/0.98

= 1.024

Step 4

= 1.024^18.25

= 50.69

Step 5

1-1.5069

= 0.5069

50.69% is the annual percentage rate for forgoing cash discount